A sample of pure CaCO3 weighing 0.3677g is dissolved in hydrochloric acid and the solution diluted to 250.0ml in a volumetric flask. A 25.00ml aliquot requires 30.26ml of an EDTA solution for titration. Calculate a) the molarity of the EDTA solution; b) the number of grams Na2H2Y•2H2O (FW = 372.2) required to prepare 500.0ml of the solution. 


1
Expert's answer
2021-01-26T03:16:14-0500

Solution:

The reaction between CaCO3 and hydrochloric acid can be described using the following equation:

CaCO3(s) + 2HCl(aq) → Ca2+(aq) + 2Cl-(aq) + H2O(l) + CO2(g)

According to the equation: Moles of CaCO3 = Moles of Ca2+


PART A:

Find the molarity of Ca2+:

(0.3677 g CaCO3)×(1 mol CaCO3/100.09 g CaCO3)×(1 mol Ca2+/1 mol CaCO3)= 0.003674 mol Ca2+ ions

Molarity of Ca2+ = Moles of Ca2+ / Volume solution = 0.003674 mol / 0.250 L = 0.01469 M Ca2+ ions

Molarity of Ca2+ = 0.01469 M Ca2+ ions


When calcium ions (Ca2+) are titrated with EDTA a relatively stable calcium complex is formed:

Ca2+ + H2Y2- = CaY2- + 2H+

According to the equation: Moles of Ca2+ = Moles of H2Y2- (EDTA)

Find the molarity of EDTA:

Moles of EDTA = (0.01469 mol Ca2+ / L) × (0.025 L) × (1 mol H2Y2- / 1 mol Ca2+) = 0.0003673 mol EDTA

Molarity of EDTA = Moles of EDTA / Volume of EDTA = 0.0003673 mol / 0.03026 L = 0.01214 M EDTA

Molarity of EDTA = 0.01214 M


PART B:

Molarity of EDTA = 0.01214 M

Molarity of EDTA = Moles of EDTA / Volume solution

Hence,

Moles of EDTA = (0.01214 mol EDTA / L) × (0.500 L / 1) = 0.00607 mol EDTA


Moles of EDTA = Mass of EDTA / Molar mass of EDTA

Hence,

Mass of EDTA = (0.00607 mol EDTA) × (372.2 g EDTA / 1 mol) = 2.2593 g

Mass of EDTA = 2.2593 g


Answer:

(a): The molarity of the EDTA solution is 0.01214 M

(b): 2.2593 grams of EDTA required